OpenStudy (anonymous):

Say I had a question like pi^x=5.4 and I wanted to find x. How would I go about this algebraically?

OpenStudy (lalaly):

ln(pi^x)=ln(5.4) xln(pi)=ln(5.4) x=ln(pi)/ln(5.4)

OpenStudy (anonymous):

Ok but is there an easy way to do ln(pi)/ln(5.4)=x without a calculator?

OpenStudy (anonymous):

@lalaly when you divide ln(pi)??? to get x xln(pi)=ln(5.4) x=ln(pi)/ln(5.4)

OpenStudy (lalaly):

sorry x=ln(5.4)/ln(pi)